OpenStudy (anonymous):

The volume of a cube is 42 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

The volume of cube \[ V= a^3 \]

OpenStudy (anonymous):

The volume of pyramid inside is \[ V_p=\frac 1 3 a^2 h=\frac 1 3 a^2 a=\frac 1 3 a^3=\frac 13( 42)=14 \]

OpenStudy (anonymous):

is that the answer @eliassaab

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

YW